Circular No. 6268

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

51 PEGASI W. B. Landsman, Hughes STX; T. Simon, University of Hawaii; M.

Mayor, Geneva Observatory; and I. Skillen, VILSPA Observatory, communicate: “We obtained an International Ultraviolet Explorer (IUE) lowdispersion, short-wavelength observation of 51 Peg on Nov. 25 (exposure time 360 min). The spectrum is that of a normal G dwarf, with no continuum detected shortward of 165.0 nm. At 160.0 nm, we place an upper limit on the continuum flux of 2.43 × 10^{−15} erg cm^{−2} s^{−1} A^{−1}. From this limit, we infer that if a normal, 0.6-M_{⊙} white dwarf exists in the system, then it must be cooler than 10 000 K (for a distance of 17.4 pc). The spectrum between 170.0 and 195.0 nm is a very close match to an archival IUE spectrum of α Cen A, scaled by the ratio of the V magnitudes. Very weak C II and C IV emission in 51 Peg is detected with integrated fluxes of 2.5 and 4.0 × 10^{−14} erg cm^{−2} s^{−1}, respectively — consistent with a solar-activity level and a rotation period of between 23 and 30 days.”

LEONID METEORS 1995 Various reports, including a summary from J. Rendtel, International

Meteor Organization, indicate that the Leonids in 1995 were evidently less pronounced than in 1994 (cf. IAUC 6109). Rendtel’s data from numerous observers delineate a rather broad plateau for maximum activity extending from Nov. 17.9 to 18.4 UT (ZHR ∼ 25–35).

SUPERNOVA 1995an IN UGC 3188 Magnitude and precise position from H. Tomsik, S. Binnewies, and P.

Riepe (Marl, Germany; 0.45-m f/5 reflector + CCD): Nov. 20.97 UT, 16.7, α = 4^{h}51^{m}48^{s}.54, δ = +8^{o}50′39′′.7 (equinox 2000.0).

CORRIGENDA

On IAUC 6264, HT Cassiopeiae, for 17.733, read 18.733. Add the negative observation 17.74, [13.0 (Schmeer). The earliest positive observation of the star’s brightening was that by W. Warraker on 17.912. This observation was immediately confirmed by G. Poyner (Birmingham) and communicated by G. M. Hurst (Basingstoke).

On IAUC 6265, α Monocerotid Meteors 1995, line 13 from foot, for K. Sárenczky read K. Sárneczky; for Vértes read Gerecse

1995 November 30 Daniel W. E. Green